Which of the following ideal aqueous solutions would be predicted to have the highest freezing point?

1.00 m Na2SO4
1.00 m RbCl
0.50 m KNO3
2.00 m methanol (CH3OH)
Answer #1

highest freezig point ... menas the point that will depress the LEAST

that is given by

dTf = Kf*m*i

since KF is constatn

then

m*i is the factors

for Na2SO4 --< 1*3 = 3

and methanol --> 2*1 = 2

then

the lowest

KNO3 = 0.5*2 = 1

the least number of "m*i" is given by KNO3